# Universal Beta Splatting

### v1

@@ Abstract
We introduce Universal Beta Splatting (UBS), a unified framework that generalizes 3D Gaussian Splatting to N-dimensional anisotropic Beta kernels for explicit radiance field rendering. Unlike fixed Gaussian primitives, Beta kernels enable controllable dependency modeling across spatial, angular, and temporal dimensions within a single representation. Our unified approach captures complex light transport effects, handles anisotropic view-dependent appearance, and models scene dynamics without requiring auxiliary networks or specific color encodings. UBS maintains backward compatibility by approximating to Gaussian Splatting as a special case, guaranteeing plug-in usability and lower performance bounds. The learned Beta parameters naturally decompose scene properties into interpretable without explicit supervision: spatial (surface vs. texture), angular (diffuse vs. specular), and temporal (static vs. dynamic). Our CUDA-accelerated implementation achieves real-time rendering while consistently outperforming existing methods across static, view-dependent, and dynamic benchmarks, establishing Beta kernels as a scalable universal primitive for radiance field rendering. Our project website is available at https://rongliu-leo.github.io/universal-beta-splatting/.
@@ 1 Introduction
Recent advances partially address these limitations through specialized designs. Deformable Beta Splatting (DBS) (Liu et al., 2025) improves geometric fidelity using spatially adaptive Beta kernels, but remains limited to 3D spatial dimensions and still requires separate Spherical Beta functions for view-dependent color. $N$-dimensional Gaussian extensions like 6DGS (Gao et al., 2025a) and 7DGS (Gao et al., 2025b) incorporate view and temporal dimensions through conditional distributions, enabling them to capture scattering effects and dynamic changes. However, these approaches remain constrained by the Gaussian kernel’s symmetric profile across dimensions—preventing independent control of spatial sharpness, angular specularity, and temporal dynamics.
@@ 2 Related Work
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) achieved real-time rendering through explicit anisotropic Gaussian primitives. Extensions address anti-aliasing (Yu et al., 2024; Yan et al., 2024), compression (Lee et al., 2024; Morgenstern et al., 2024), and semantic applications (Shi et al., 2024; Zhou et al., 2024; Peng et al., 2025; Xiong et al., 2025). However, 3DGS’s spherical harmonic encoding struggles with complex view-dependent effects. To better capture view dependence, $N$-dimensional Gaussians (N-DG) (Diolatzis et al., 2024) extended to higher dimensions, refined by 6D Gaussian Splatting (6DGS) (Gao et al., 2025a) using conditional slicing. While elegant, Gaussian distributions cannot produce sharp angular cutoffs without many small primitives, limiting efficiency for specular materials.
Alternative Kernel Designs. Recent work explores kernels beyond Gaussians: GES (Hamdi et al., 2024) uses generalized exponentials for edge representation; Deformable Beta Splatting (DBS) (Liu et al., 2025) introduces adaptive Beta kernels with learnable shape parameters; 3D Convex Splatting (Held et al., 2025) handles flat surfaces; 3D Student Splatting (Zhu et al., 2025) leverages heavier-tailed distributions; and Gabor Splats (Zhou et al., 2025) captures high-frequency details. While these improve specific aspects, they remain confined to 3D spatial modeling without addressing view-dependent or temporal dimensions.
@@ 3 Preliminary
3D Gaussian Splatting (3DGS). 3DGS (Kerbl et al., 2023) represents scenes as collections of anisotropic 3D Gaussians parameterized by position $\bm{\mu}\in\mathbb{R}^{3}$, opacity $\alpha$, rotation quaternion $\bm{q}$, scale $\bm{s}$, and color features $\bm{f}$. The covariance matrix $\bm{\Sigma}=\bm{R}\bm{S}\bm{S}^{\top}\bm{R}^{\top}$ is constructed from rotation and scale matrices. Rendering uses alpha-composited 2D projections:
$\bm{C}(\bm{x})=\sum_{i=1}^{N}c_{i}\alpha_{i}\mathcal{G}\big(r_{i}(\bm{x})\big)\prod_{j=1}^{i-1}\Big(1-\alpha_{j}\mathcal{G}\big(r_{j}(\bm{x})\big)\Big),$ (1)
where $\mathcal{G}(r)$ is the 2D Gaussian and $r_{i}(\bm{x})$ is the Mahalanobis distance.
Deformable Beta Splatting (DBS). DBS (Liu et al., 2025) replaces Gaussians with adaptive Beta kernels:
$\mathcal{B}(x;b)=(1-x)^{\beta(b)},\quad\beta(b)=4\exp(b),\quad x\in[0,1],$ (2)
where $b$ controls kernel shape (negative: flat surfaces, positive: peaked details, zero: Gaussian-like).
N-Dimensional Gaussian. 6DGS (Gao et al., 2025a) and 7DGS (Gao et al., 2025b) extend 3DGS through conditional distributions. 6DGS incorporates viewing direction into a 6D joint distribution:
@@
Conditioning on view $d$ yields conditional covariance $\bm{\Sigma}_{\text{cond}}=\bm{\Sigma}_{p}-\bm{\Sigma}_{pd}\bm{\Sigma}_{d}^{-1}\bm{\Sigma}_{pd}^{\top}$, conditional mean $\bm{\mu}_{\text{cond}}=\bm{\mu}_{p}+\bm{\Sigma}_{pd}\bm{\Sigma}_{d}^{-1}(d-\bm{\mu}_{d})$, and modulated opacity $\alpha_{\text{cond}}=\alpha\cdot\exp(-\lambda(d-\bm{\mu}_{d})^{\top}\bm{\Sigma}_{d}^{-1}(d-\bm{\mu}_{d}))$.
@@ 4 Method
Figure 2: Visualization of decomposition. Our learned Beta parameters provide interpretable scene decomposition across both spatial, angular, and temporal dimension without explicit supervision.
@@ 4.1 N-Dimensional Beta Kernel
The core innovation of UBS is the $N$-dimensional anisotropic Beta kernel that models the universal radiance function:
$\sigma(\bm{x},\bm{q})=B\bigl(\bm{x},\bm{q};\,\bm{\mu},\,\bm{\Sigma},\,\bm{b}\bigr),$ (4)
where $\bm{x}\in\mathbb{R}^{3}$ represents spatial coordinates, the query $\bm{q}\in\mathbb{R}^{N-3}$ encodes additional dimensions (view direction, time, etc.), $\bm{\mu}\in\mathbb{R}^{N}$ is the mean vector, $\bm{\Sigma}\in\mathbb{R}^{N\times N}$ is the covariance matrix, and $\bm{b}\in\mathbb{R}^{N-2}$ controls the Beta shape parameters across dimensions (Betas are shared in spatial domain). The shape parameters $\bm{b}$ are transformed to ensure positive Beta exponents: $\beta_{i}=4\exp(b_{i})$, enabling each dimension to adopt optimal kernel forms—from flat (negative $b_{i}$) to peaked (positive $b_{i}$).
In addition to the primitive shape, each kernel primitive also carries an opacity $o\in[0,1]$ and an RGB color $\bm{c}\in[0,1]^{3}$. Crucially, because the Beta kernel can natively incorporate view and temporal dimensions through $\bm{q}$, no auxiliary color encoding is required to model view-dependent appearance or dynamics. This unified representation significantly reduces the per-primitive parameter count: for static scenes, UBS achieves a 41% parameter reduction compared to 3DGS, and for dynamic scenes, UBS uses 73% fewer parameters than 4DGS, as detailed in Table 5 in the Appendix. In contrast, existing methods rely on auxiliary encodings: 3DGS, 6DGS, and 7DGS employ 48-parameter spherical harmonics, 4DGS requires 144-parameter 4D spherical harmonics, and DBS needs additional spherical Beta functions for color modeling.
@@ 4.2 Spatial-Orthogonal Cholesky Parameterization
$\bm{L}=\begin{pmatrix}\bm{R}_{x}\,\text{diag}(\bm{s}_{x})&\bm{0}\\ \bm{L}_{qx}&\text{diag}(\bm{s}_{q})\end{pmatrix}$, (6)
where $\bm{R}_{x}\in\text{SO}(3)$ is the spatial rotation computed via first-order Taylor approximation $\bm{R}_{x}\approx\bm{I}+\bm{A}_{x}$ of the matrix exponential (with $\bm{A}_{x}$ skew-symmetric), $\bm{s}_{x}$ are spatial scales, $\bm{s}_{q}$ are scales for other dimensions, and $\bm{L}_{qx}$ encodes cross-correlations. As shown in Table 6 (Appendix), this first-order approximation provides nearly identical fidelity while greatly reducing computation. The covariance $\bm{\Sigma}=\bm{L}\bm{L}^{\top}$ maintains explicit 3D geometric structure while allowing expressive correlations across view and time.
@@ 4.3 Beta-Modulated Conditional Slicing
$\bm{\mu}_{x\mid q}$ $=\bm{\mu}_{x}+\bm{\Sigma}_{xq}\,\bm{\Sigma}_{q}^{-1}\,\text{Diag}(\bm{\beta}_{q})\,(\bm{q}-\bm{\mu}_{q}),$ (7)
$\bm{\Sigma}_{x\mid q}$ $=\bm{\Sigma}_{x}-\bm{\Sigma}_{xq}\,\bm{\Sigma}_{q}^{-1}\,\text{Diag}(\bm{\beta}_{q})\,\bm{\Sigma}_{qx},$ (8)
where $\text{Diag}(\bm{\beta}_{q})=\text{Diag}([\beta_{q_{1}},\ldots,\beta_{q_{C}}])$ applies dimension-specific Beta modulation to the non-spatial conditioning terms while preserving isotropic spatial structure.
Product-Form Opacity Gate. We adopt a factorized opacity with inherent multidimensional support:
$\bm{d}^{\text{raw}}=\bm{\Sigma}_{q}^{-1}(\bm{q}-\bm{\mu}_{q}),\quad\bm{d}$ $=\max\big(\tanh{(\frac{\bm{d}^{\text{raw}}}{2})},\;0\big)\in[0,1)^{C},\quad o(\bm{q})=o\prod_{i=1}^{C}(1-d_{i})^{4\beta_{q_{i}}},$ (9)
where $\tanh$ and $\max$ transformations ensure bounded distance measures $\bm{d}\in[0,1)^{C}$, satisfying Beta kernel input definition. The product form allows independent control over each non-spatial dimension’s contribution.
@@
$\sigma(\bm{x},\bm{q})=B\!\left(\bm{x};\,\bm{\mu}_{x\mid q},\,\bm{\Sigma}_{x\mid q},\,b_{x}\right)\cdot o(\bm{q}),$ (10)
@@ 4.4 Universal Compatibility and Interpretability
Universal Compatibility. UBS provides backward compatibility by design, reducing to existing methods under specific parameter settings. For $N=3$ (pure spatial), UBS reduces to Deformable Beta Splatting with $\sigma(\bm{x})=B(\bm{x};\bm{\mu}_{x},\bm{\Sigma}_{x},b_{x})\cdot o$, and further approximates to standard 3DGS when $b_{x}=0$ (Gaussian limit). For higher dimensions with all Beta parameters zero ($\bm{b}=\bm{0}$), UBS approximates corresponding Gaussian methods: 6DGS for view-dependent rendering ($N=6$) and 7DGS for dynamic scenes ($N=7$). This guaranteed compatibility ensures UBS serves as a seamless drop-in replacement while providing performance lower bounds.
@@ 4.5 Optimization and Implementation
For primitive management, we adopt the kernel-agnostic MCMC optimization from DBS (Liu et al., 2025), which provides stable, distribution-preserving updates independent of kernel form.

### v2

@@ Abstract
We introduce Universal Beta Splatting (UBS), a unified framework that generalizes 3D Gaussian Splatting to N-dimensional anisotropic Beta kernels for explicit radiance field rendering. Unlike fixed Gaussian primitives, Beta kernels enable controllable dependency modeling across spatial, angular, and temporal dimensions within a single representation. Our unified approach captures complex light transport effects, handles anisotropic view-dependent appearance, and models scene dynamics without requiring auxiliary networks or specific color encodings. UBS maintains backward compatibility by approximating to Gaussian Splatting as a special case, guaranteeing plug-in usability and lower performance bounds. The learned Beta parameters naturally decompose scene properties without explicit supervision: spatial (surface vs. texture), angular (diffuse vs. specular), and temporal (static vs. dynamic). Our CUDA-accelerated implementation achieves real-time rendering while consistently outperforming existing methods across static, view-dependent, and dynamic benchmarks, establishing Beta kernels as a scalable universal primitive for radiance field rendering. Our project website is available at https://rongliu-leo.github.io/universal-beta-splatting/.
@@ 1 Introduction
Recent advances partially address these limitations through specialized designs. Deformable Beta Splatting (DBS) (Liu et al., 2025a) improves geometric fidelity using spatially adaptive Beta kernels, but remains limited to 3D spatial dimensions and still requires separate Spherical Beta functions for view-dependent color. $N$-dimensional Gaussian extensions like 6DGS (Gao et al., 2025a) and 7DGS (Gao et al., 2025b) incorporate view and temporal dimensions through conditional distributions, enabling them to capture scattering effects and dynamic changes. However, these approaches remain constrained by the Gaussian kernel’s symmetric profile across dimensions—preventing independent control of spatial sharpness, angular specularity, and temporal dynamics.
@@ 2 Related Work
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) achieved real-time rendering through explicit anisotropic Gaussian primitives. Extensions address anti-aliasing (Yu et al., 2024; Yan et al., 2024), compression (Lee et al., 2024a; Morgenstern et al., 2024), and semantic applications (Shi et al., 2024; Zhou et al., 2024; Peng et al., 2025; Xiong et al., 2026). However, 3DGS’s spherical harmonic encoding struggles with complex view-dependent effects. To better capture view dependence, $N$-dimensional Gaussians (N-DG) (Diolatzis et al., 2024) extended to higher dimensions, refined by 6D Gaussian Splatting (6DGS) (Gao et al., 2025a) using conditional slicing. While elegant, Gaussian distributions cannot produce sharp angular cutoffs without many small primitives, limiting efficiency for specular materials.
Alternative Kernel Designs. Recent work explores kernels beyond standard 3D Gaussians. Several methods modify the kernel shape to better capture geometric structure: GES (Hamdi et al., 2024), TNT-GS (Liu et al., 2025b), Quadratic Gaussian Splatting (Zhang et al., 2025), 3D Half-Gaussian Splatting (Li et al., 2025), and Disc-GS (Qu et al., 2024) propose generalized exponentials, truncated, second-order, half, and discontinuity-aware Gaussian splats. Other approaches expand the distribution family, including Deformable Beta Splatting (DBS) (Liu et al., 2025a), 3D Student Splatting (Zhu et al., 2025), and Gabor Splats (Zhou et al., 2025) for sharper edges or heavier-tailed behavior. For appearance modeling, Textured Gaussians (Chao et al., 2025) enrich per-splat color detail. Beyond volumetric kernels, 3D Convex Splatting (Held et al., 2025b), Triangle Splatting (Held et al., 2025a), and Deformable Radial Kernel Splatting (Huang et al., 2025) introduce other geometric primitives as alternatives to ellipsoids. While these methods improve specific aspects of spatial reconstruction, they remain confined to 3D spatial kernels, without extending to unified high-dimensional splatting that jointly models spatial, angular, or temporal dimensions.
@@ 3 Preliminary
3D Gaussian Splatting (3DGS). 3DGS (Kerbl et al., 2023) represents scenes as collections of anisotropic 3D Gaussians parameterized by position $\bm{\mu}\in\mathbb{R}^{3}$, opacity $o$, rotation quaternion $\bm{q}$, scale $\bm{s}$, and color features $\bm{f}$. The covariance matrix $\bm{\Sigma}=\bm{R}\bm{S}\bm{S}^{\top}\bm{R}^{\top}$ is constructed from rotation and scale matrices. To render an image, each 3D Gaussian ellipsoid is projected into the 2D image plane using the viewing transformation $\bm{W}$. This yields a 2D mean $\bm{\mu}^{\prime}\in\mathbb{R}^{2}$ and a 2D covariance matrix $\bm{\Sigma}^{\prime}=\bm{JW\Sigma W^{\top}J^{\top}}$, where $\bm{J}$ is the Jacobian of the projection. For a pixel location $\bm{x}\in\mathbb{R}^{2}$, the Mahalanobis distance to the projected splat is
$r_{i}(\bm{x})=(\bm{x}-\bm{\mu}_{i}^{\prime})^{\top}\bm{\Sigma}_{i}^{\prime^{-1}}(\bm{x}-\bm{\mu}_{i}^{\prime}).$ (1)
The Gaussian contribution to the pixel is then computed using a Gaussian kernel multiplied by opacity:
$\sigma(\bm{x})=\mathcal{G}(\bm{x};\bm{\mu},\bm{\Sigma})\cdot o=e^{-\frac{1}{2}r_{i}(\bm{x})}\cdot o.$ (2)
Finally, after sorting the primitives in front-to-back order, the pixel color is obtained using standard alpha compositing:
$\bm{C}(\bm{x})=\sum_{i=1}^{N}c_{i}\sigma_{i}(\bm{x})\prod_{j=1}^{i-1}(1-\sigma_{j}(\bm{x})).$ (3)
Deformable Beta Splatting (DBS). DBS (Liu et al., 2025a) replaces Gaussian kernels with adaptive and bounded Beta kernels, defining density as $\sigma(\bm{x})=\mathcal{B}(\bm{x};\bm{\mu},\bm{\Sigma},b)\cdot o$, where
$\mathcal{B}(\bm{x};\bm{\mu},\bm{\Sigma},b)=(1-r(\bm{x}))^{\beta(b)},\quad\beta(b)=4\exp(b),\quad b\in\mathbb{R},\quad r(\bm{x})\in[0,1).$ (4)
The Beta kernel closely approximates a scaled Gaussian when initialization. As training progresses, the learnable parameter $b$ allows the kernel shape to adapt: negative $b$ produces flatter profiles with sharper boundaries suitable for solid geometry, while positive $b$ yields sharper peaks that better capture high-frequency details. This adaptability enables DBS to model diverse geometric structures more effectively than fixed Gaussian kernels. Appendix D includes more background and motivation.
N-Dimensional Gaussian. 6DGS (Gao et al., 2025a) and 7DGS (Gao et al., 2025b) extend 3DGS through conditional distributions. The key idea is to model a joint distribution over spatial and non-spatial dimensions (view/time), then condition on the query to obtain a view-specific or time-specific 3D Gaussian for rendering. 6DGS incorporates viewing direction into a 6D joint distribution:
@@
Conditioning on view $d$ yields conditional covariance $\bm{\Sigma}_{\text{cond}}=\bm{\Sigma}_{p}-\bm{\Sigma}_{pd}\bm{\Sigma}_{d}^{-1}\bm{\Sigma}_{pd}^{\top}$, conditional mean $\bm{\mu}_{\text{cond}}=\bm{\mu}_{p}+\bm{\Sigma}_{pd}\bm{\Sigma}_{d}^{-1}(d-\bm{\mu}_{d})$, and modulated opacity $o_{\text{cond}}=o\cdot\exp(-\lambda(d-\bm{\mu}_{d})^{\top}\bm{\Sigma}_{d}^{-1}(d-\bm{\mu}_{d}))$.
@@ 4 Method
Figure 2: Visualization of decomposition. Our learned Beta parameters provide interpretable scene decomposition across spatial, angular, and temporal dimensions without explicit supervision.
@@ 4.1 N-Dimensional Beta Kernel
The core innovation of UBS is the $N$-dimensional anisotropic Beta kernel that generalizes the DBS density function to higher dimensions:
$\sigma(\bm{x},\bm{q})=\mathcal{B}\bigl(\bm{x},\bm{q};\,\bm{\mu},\,\bm{\Sigma},\,\bm{b}\bigr)\cdot o,$ (6)
where $\bm{x}\in\mathbb{R}^{3}$ represents spatial coordinates, the query $\bm{q}\in\mathbb{R}^{N-3}$ encodes additional dimensions (view direction, time, etc.), $\bm{\mu}\in\mathbb{R}^{N}$ is the mean vector, $\bm{\Sigma}\in\mathbb{R}^{N\times N}$ is the covariance matrix, $\bm{b}\in\mathbb{R}^{N-2}$ controls the Beta shape parameters across dimensions (Betas are shared in spatial domain), and $o\in[0,1]$ is the opacity. The shape parameters $\bm{b}$ are transformed to ensure positive Beta exponents: $\beta_{i}=4\exp(b_{i})$, enabling each dimension to adopt optimal kernel forms—from flat (negative $b_{i}$) to peaked (positive $b_{i}$).
In addition to the primitive shape, each kernel primitive also carries an RGB color $\bm{c}\in[0,1]^{3}$. Crucially, because the Beta kernel can natively incorporate view and temporal dimensions through $\bm{q}$, no auxiliary color encoding is required to model view-dependent appearance or dynamics. This unified representation significantly reduces the per-primitive parameter count: for static scenes, UBS achieves a 41% parameter reduction compared to 3DGS, and for dynamic scenes, UBS uses 73% fewer parameters than 4DGS, as detailed in Table 7 in the Appendix B. In contrast, existing methods rely on auxiliary encodings: 3DGS, 6DGS, and 7DGS employ 48-parameter spherical harmonics, 4DGS requires 144-parameter 4D spherical harmonics, and DBS needs additional spherical Beta functions for color modeling.
@@ 4.2 Spatial-Orthogonal Cholesky Parameterization
$\bm{L}=\begin{pmatrix}\bm{R}_{x}\,\text{diag}(\bm{s}_{x})&\bm{0}\\ \bm{L}_{qx}&\bm{L}_{q}\end{pmatrix}$, (8)
where $\bm{R}_{x}\in\text{SO}(3)$ is the spatial rotation computed via first-order Taylor approximation $\bm{R}_{x}\approx\bm{I}+\bm{A}_{x}$ of the matrix exponential (with $\bm{A}_{x}$ skew-symmetric), $\bm{s}_{x}$ are spatial scales, $\bm{s}_{q}$ are scales for other dimensions, and $\bm{L}_{qx}$ encodes cross-correlations. As shown in Table 10 (Appendix C), this first-order approximation provides nearly identical fidelity while greatly reducing computation. The covariance $\bm{\Sigma}=\bm{L}\bm{L}^{\top}$ maintains explicit 3D geometric structure while allowing expressive correlations across view and time. More theoretical and empirical analysis are available in Appendix C.
@@ 4.3 Beta-Modulated Conditional Slicing
$\bm{\mu}_{x\mid q}$ $=\bm{\mu}_{x}+\bm{\Sigma}_{xq}\,\bm{\Sigma}_{q}^{-1}\,\text{diag}(\bm{\tilde{\beta}}_{q})\,(\bm{q}-\bm{\mu}_{q}),$ (9)
$\bm{\Sigma}_{x\mid q}$ $=\bm{\Sigma}_{x}-\bm{\Sigma}_{xq}\,\bm{\Sigma}_{q}^{-1}\,\text{diag}(\bm{\tilde{\beta}}_{q})\,\bm{\Sigma}_{qx},$ (10)
where $\text{diag}(\bm{\tilde{\beta}}_{q})=\text{diag}([\beta_{q_{1}},\ldots,\beta_{q_{C}}]_{\leq 1})$ applies dimension-specific Beta modulation to the non-spatial conditioning terms while preserving isotropic spatial structure.
Beta-Modulated Opacity. We adopt a factorized opacity with inherent multidimensional support:
$d_{i}^{\text{raw}}=(\hat{\bm{L}}_{q}^{-1}(\bm{q}-\bm{\mu}_{q}))^{2}_{i},\quad d_{i}$ $=\tanh(d_{i}^{\text{raw}})\in[0,1),\quad o(\bm{q})=o\prod_{i=1}^{C}(1-d_{i})^{4\beta_{q_{i}}},$ (11)
where $\hat{\bm{L}}_{q}$ denotes the Cholesky factor of $\bm{\Sigma}_{q}$, such that $\hat{\bm{L}}_{q}^{-1}\hat{\bm{L}}_{q}^{-\top}=\bm{\Sigma}_{q}^{-1}$. Then $\tanh$ maps per-dimension Mahalanobis distance $d_{i}^{\text{raw}}$ into a bounded value $d_{i}\in[0,1)$, satisfying the Beta kernel input definition. The modulated product form allows independent control over each non-spatial dimension’s contribution.
@@
$\sigma(\bm{x},\bm{q})=\mathcal{B}\!\left(\bm{x};\,\bm{\mu}_{x\mid q},\,\bm{\Sigma}_{x\mid q},\,b_{x}\right)\cdot o(\bm{q}),$ (12)
@@ 4.4 Universal Compatibility and Interpretability
Universal Compatibility. UBS provides backward compatibility by design, reducing to existing methods under specific parameter settings. For $N=3$ (pure spatial), UBS reduces to Deformable Beta Splatting with $\sigma(\bm{x})=\mathcal{B}(\bm{x};\bm{\mu}_{x},\bm{\Sigma}_{x},b_{x})\cdot o$, and further approximates to standard 3DGS when $b_{x}=0$ (Gaussian limit). For higher dimensions with all Beta parameters zero ($\bm{b}=\bm{0}$), UBS approximates corresponding Gaussian methods: 6DGS for view-dependent rendering ($N=6$) and 7DGS for dynamic scenes ($N=7$). This guaranteed compatibility ensures UBS serves as a seamless drop-in replacement while providing performance lower bounds.
@@ 4.5 Optimization and Implementation
For primitive management, we adopt the kernel-agnostic MCMC optimization from DBS (Liu et al., 2025a), which provides stable, distribution-preserving updates independent of kernel form.